OpenStudy (anonymous):

how do I know when trying to solve a quadractic equation and having to graph it, if it opens open or down on the graph?

OpenStudy (anonymous):

if the term in x^2 has a positive coefficient then it opens up on the graph - if its negative it opens down

OpenStudy (anonymous):

This is the equation -0.2x^2+12x+11=0

OpenStudy (anonymous):

crudely speaking x^2 is like a U and -x^2 is like an upside down U !

OpenStudy (anonymous):

ok thanks!!!!

OpenStudy (anonymous):

micaiah graph would open down because of the -0.2x^2

OpenStudy (anonymous):

Could you help me solve this? x=1on that same equation.

OpenStudy (anonymous):

thats u of course!

OpenStudy (anonymous):

oh - you want the points of intersection of the quadratic and x=-1 right?

OpenStudy (anonymous):

Well let me explain the problem to you. Suppose you are an event coordinator. The box office manager wants a projected ticket sales report. The quadratic equation that I gave models the expected number of ticket sales for each day (x). (x=1 is the day the tickets go on sale)

OpenStudy (anonymous):

On day one, sales projected = 0.2 + 11 + 12 = 23.2 - i suppose that becomes 23 as tickets sold must be whole numbers. Day 2 sales will be found by plugging in x=2 and so on

OpenStudy (anonymous):

you can also find when the maximum number of sales occur by finding the vertex (peak) of the curve

OpenStudy (anonymous):

this -b/2a for the function ax^2 + bx + c for our quadratic it is -12 / 2(-0.2) = 30 ie on 30th day

OpenStudy (anonymous):

on that day number of sales = -0.2 *30^2 +12(30) + 11 tickets = 191

OpenStudy (anonymous):

sorry - i calculated day one incorrectly - its -0.2 + 11+12 tickets = 22.8 - we can call that 23 tickets

OpenStudy (anonymous):

30 days corresponds to the peak of the curve - after that day tickets sales will decrease

OpenStudy (anonymous):

if u drew the graph you can read off the projected sales for any day. graph will look like a symmetrical arch. The math name for it is a parabola

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

Thank youu so much. I really needed your help. I really appreciate it.